media%2F823%2F82313742-08f2-4149-a84f-5e
Answer #1

From the figure: Depth - (18+18-y) pg(36-y)32dy 1000 kg/m (9.8 m/s) (32)(36-y)dy = ( F313600(36-y)d 313600| 36-r = 313600(36-